OpenStudy (anonymous):

what are the solutions of |x+6|>=5? (this is equal to or greater than)

OpenStudy (anonymous):

absolute value is removed by dividing by positive or negative \[x >= -1\] or \[x <= -11\]

OpenStudy (anonymous):

\[\left\{ x|x \le -11 \right\}\cup \left\{ x|x \ge -1 \right\} or \left\{ x|x \le -11 \right\}\cup \left\{ x|x \ge -1 \right\} or \left\{ x|x \le-1\left\{ x|x \ge-11 \right\} \right\} or \left\{ x|x \le-1 \right\}\left\{ x|x \ge -11 \right\}\]